OpenStudy (anonymous):

simplify: 5x-(7x+7y)

OpenStudy (kropot72):

First remove the brackets. To do this, multiply each term inside the brackets by -1. Please post the result.

OpenStudy (anonymous):

So it is 5x-7x+7y

OpenStudy (kropot72):

Not quite. What is \[7y \times -1\]

OpenStudy (anonymous):

7xy-1?

OpenStudy (anonymous):

oh is that a multiplication sign?

OpenStudy (kropot72):

Not really. If you multiply 7y by negative 1 the result is negative 7y: \[7y \times-1=-7y\]

OpenStudy (anonymous):

ok that makes sense

OpenStudy (kropot72):

So when the brackets are removed the result is: \[5x-(7x+7y)=5x-7x-7y\] Now you need to add the terms in x, taking care to observe the signs. Can you do that?

OpenStudy (anonymous):

ok so then you would get -2x-7y

OpenStudy (kropot72):

Correct! Good work.

OpenStudy (anonymous):

Thank you :)

OpenStudy (kropot72):

You're welcome :)